Dan is considering the purchase of Super Technology, Inc. bonds that were issued 6 years ago. When the bonds were originally sold they had a 24-year maturity and a 8.25 percent coupon interest rate, paid annually. The bond is currently selling for $883. Par value of the bond is $1,000. What is the yield to maturity on the bonds if you purchased the bond today?
Answer #1

To find the YTM, we need to put the following values in the financial calculator:

INPUT 24-6=18 -883 8.25%x1,000=82.50 1,000
TVM N I/Y PV PMT FV
OUTPUT 9.64

So, YTM = 9.64%